OpenStudy (idealist10):

How to integrate lnabs(e^(x)+1)/e^x dx?

OpenStudy (idealist10):

\[\frac{ \ln \left| e^x+1 \right| }{ e^x }dx\]

OpenStudy (idealist10):

@hartnn @phi

hartnn (hartnn):

how about \(u = e^{-x}\) ? tried?

hartnn (hartnn):

it works out :) try it! and let me know if u get stuck

OpenStudy (idealist10):

You mean the u-substitution?

hartnn (hartnn):

yes

hartnn (hartnn):

\(e^x+1 = \dfrac{1+e^{-x}}{e^{-x}}\)

OpenStudy (idealist10):

How so?

hartnn (hartnn):

\(e^x+1 = \dfrac{(e^x+1)e^{-x}}{e^{-x}}= \dfrac{1+e^{-x}}{e^{-x}}\)

OpenStudy (idealist10):

But how does that help us to find the integral? u=e^-x du=-e^-x dx And?

hartnn (hartnn):

\(-\int \ln (\dfrac{u+1}{u})du\) got this? ln A/B = ln A-ln B

hartnn (hartnn):

i am assuming we can use the standard integral of ln x

hartnn (hartnn):

\(\int \ln x dx = x (\ln x -1)+c\)

hartnn (hartnn):

just use that to get these : \(\int \ln u dx -\int \ln (u+1) dx\)

hartnn (hartnn):

if we cannot use it directly, we will need uv rule

OpenStudy (idealist10):

Shouldn't we use the method of integration by parts in the first place?

hartnn (hartnn):

trying that now... \(\ln |e^x+1| (-e^{-x}) -\int \dfrac{e^x}{e^x+1}\times e^{-x}dx \) yeah, works out...didn't find it shorter though

hartnn (hartnn):

if we could directly use the result for ln x, then the u subs is easier/shorter. otherwise, by parts from the beginning is easier.

OpenStudy (idealist10):

Well, the result I got is \[-\ln \left| e^x+1 \right|e ^{-x}+x-\ln \left| e^x+1 \right|+C\]

hartnn (hartnn):

\(\ln |e^x+1| (-e^{-x}) -\int \dfrac{e^{-x}}{e^{-x}+1}dx \\ u = 1+e^{-x}\)

hartnn (hartnn):

\(\ln |e^x+1| (-e^{-x}) -\int \dfrac{e^{-x}}{e^{-x}+1}dx \\ u = 1+e^{-x} \\ \ln |e^x+1| (-e^{-x}) +\int \dfrac{1}{u}du \\\ln |e^x+1| (-e^{-x}) +\ln (e^{-x}+1)+c \)

hartnn (hartnn):

did i make a sign mistake anywhere?

OpenStudy (idealist10):

No.

hartnn (hartnn):

aah got it! \(\ln |e^x+1| (-e^{-x}) -\int \dfrac{e^x}{e^x+1}\times (-e^{-x})dx\)

hartnn (hartnn):

\(\ln |e^x+1| (-e^{-x}) +\int \dfrac{e^{-x}}{e^{-x}+1}dx \\ u = 1+e^{-x} \\ ...\)